In 3 years, Jeff will be three times as old as his sister iya. A year ago, Jeff has 7 times as add as iya. How old are they now ?
Answer by Cromlix(4381)

Hi there,
In 3 years, Jeff will be three
times as old as his sister iya.
J + 3 = 3(I + 3)
J + 3 = 3I + 9
J = 3I + 6.......(1)
A year ago, Jeff has 7 times
as old as iya.
J - 1 = 7(I - 1).....(2)
Replace J in Eq(2) with
3I + 6 from Eq(1)
J - 1 = 7(I - 1)
3I + 6 - 1 = 7(I - 1)
3I + 5 = 7I - 7
Collect like terms
3I - 7I = - 7 - 5
- 4I = - 12
I = -12/-4
I = 3
Jeff is 15 years old
Iya is 3 years old
Hope this helps :-)